what is 6% of 1.8 million over 10 years. what would be monthly for 10 years?

what is 6% of 1.8 million ?
Answer #1

Using the concept of Simple Interest:
Principal (P) = 1.8 million
Rate (R) = 6%
Time (T) = 10 years
Therefore, Simple Interest = P * R * T = 1.8mill * 6% * 10 = 1.08 million
That is 1.08 million for 10 years, meaning 0.108 million per year

While computing monthly, total number of months = 120
Therefore, total simple interest divided by total number of months = 1.08 million / 120 = 9000 per month

6% of 1.8 million = 108,000

Using the concept of compound interest
Amount (A) = P * ( 1 + R / n ) ^ nT where n = number of times rate is to be applied

therefore, A = 1.8mill * ( 1 + 0.06 / 1 ) ^ 10 = 3,223,525.85
Therefore, Total interest = A - P = 3,223,525.85 - 1800000 = 1,423,525.85
That is, average annual interest = 1,423,525.85 / 10 = 142,352.585

While computing monthly, n= 12

Therefore, A = 1.8mill * [ 1 + (0.06 / 12) ] ^ (12*10) = 3,274,914.12
Therefore, Total interest = A - P = 3,274,914.12 - 1800000 = 1,474,914.12
That is, average monthly interest = 1,474,914.12 / 120 = 12,290.951

6% of 1.8 million = 108,000